PartnerObserving Log for 2007 May 11/12

Technical Observers: J. Hill, A. DiPaola, R. Speziali
Telescope Operator: E. Lundin
Instrument: LBC-Blue

After waiting for clouds/virga to pass, open dome at 03:10.

Testing new camera software, even in ways that we did not anticipate.

03:34 - 1sec B exposure of pointing star

Image did not arrive at /Repository. LBC archive machine lbcarchive2=lbcdh1 was hung. Apparently it hung during Andrea's 100 bias frame test this afternoon. Rebooted it from the reset button. That brought the archive machine to life, and it tftp'd the images from the Windows machine. However, it has failed to mount the NFS2 disk "telescope" aka "Repository". Rebooted again from the command line at 03:50 UT. No help. Images remain in /data/store/indir/. I ftp'd the recent image to /tmp on lbtmu03. But ds9 was not able to open them as a mosaic. At 04:00:09 (a cron task starts?) images from Andrea's bias tests start appearing in /Repository. At first they were arriving about every 10 sec, but as time goes on the arrival rate has slowed to one every 3.5 minutes. This is the arrival rate of images appearing in /Repository/.

While waiting for that, it seems that images written today cannot have the header for extension 1 read. ds9 says it cannot display the image. IRAF imheader gets a segmentation fault on extension1. IDL says: Processing: /tmp/lbcb.20070512.034437.fits
Invalid extension header encountered
XTENSION keyword missing
Only 0 extensions processed
FITS_READ ERROR: EOF encountered

Images between 21:00 UT on May 11 and 04:30 UT on May 12 are corrupt and should not be used. Problem was corrected with images starting at 04:50.

04:52 - 1sec B exposure of pointing star (default focus) with all global offsets zeroed

04:58 - 1sec B exposure of pointing star (default focus) with RX=+30 (moves the star up on the LBC-Blue array)

05:06 - Vfastextra to focus at the position of Lockman Block 4 = 10:52 +57:21 Az=327 El=56 Adjusted collimation with dRX=+3 arcsec and dY=+0.134 mm, plus dRY=+10 and dX=-0.446 to get the collimation back to the pointing model. But, this seems to have made the collimation worse.......051527 ..... so reverse dY=-0.134 ....052205

05:28 - Egami_Lockman_VLA_U_3 - Clouds are thinning, but it certainly is not photometric. Guiding failed on Dither 1. Manually added -300 nm of Z4 to adjust focus during dither 2. Manually added another -150 nm of Z4 during dither 3. It is clear that the focus offset of SDT_Uspec is not as good as that of U-BESSEL. Guiding failed on dither 5 although it says "working" with 32 sec exposure.

Adjusted the SDT_Uspec focus offset from 0.486 to 0.450 mm in leftchannelfocus.dat

06:08 - Clear active optics corrections and refocus from scratch with Vfastextra

06:17 - Had to reboot the LBC Archive machine again........Now it is running at full speed!

06:28 - There was a big jump (200 pixels) in image position during the extrafocal image. So might not be the best possible AO correction.

06:33 - Egami_Lockman_VLA_U_4 - Tracking/Guiding/Seeing are quite variable during dither 1. Guider (correctly) reports image sizes from 1.1 to 1.9 arcsec, but the image turns out at 1.5 arcsec. Guiding fails on dither 2. Highly variable seeing with guiding continues through dithers 3, 4, 5. Guiding failed on dither 5 in ~2 arcsec seeing.

07:13 - Vfastextra at the position of Pizzella NGC5719

07:17 - apizzella_1 - seeing varies 1.4 to 1.9 arcsec during dither 1. Improving slightly in the next dithers.

07:40 - Vfastextra

07:43 - apizzella_2 - variable seeing continues. One of the images has a jump/tail on the image.

08:06 - Vfastextra

08:08 - apizzella_1 (second observation) -

08:40 - apizzella_2 (second observation) - thin clouds drifting by, variable seeing ~1.5-1.9 arcsec in U at 1.4 airmasses.

Out of SDT targets, so doing other stuff..............and listening to the LBT LGS meeting on video.

09:08 - Vfastextra at UGC10214

09:12 - UGC_10214_U - 10 dither x 164 sec, seeing improving a little: U images varying varying between 1.2 and 1.5 arcsec at 1.1 airmasses Manual focus adjustment -200 nm Z4 during dither 7.

With improved seeing, we go back to a Zaritsky SDT field.....

09:51 - Vfastextra at the position of NGC_5585

09:55 - NGC_5585_U, 10-step, but now at higher airmass we have 1.2 - 1.5 arcsec images. Guiding fails on dither 8 in ~2 arcsec seeing.

10:33 - Vfastextra at El=41

10:36 - NGC_5585_V, 10-step - It's a pretty galaxy, but we've got 1.3-1.6 arcsec images through thin clouds at 1.5 airmasses.

11:19 - 1sec B exposure of pointing star (985, 2945) Az=97 El=59 11:24 - 1sec B exposure of pointing star (955, 2940) Az=98 El=21 11:30 - 1sec B exposure of pointing star (968, 2909) Az=158 El=38 11:33 - same star with azoffset=+120 arcsec (1366, 2781) I corrected the collimation model, but active optics must be overriding it.

Thin strips of cirrus in the sky during these flats.

11:40 - SkyFlatG_5 x10 Az=10 El=72 40000, 45000, 55000, saturated......

11:45 - SkyFlatUspec_5 x8 10000, ..............21000

11:51 - SkyFlatUspecRot180_5 x8 32000, 39000, 49000, saturated...............

Close 12:00 T=9.5degC D=-6degC Wind 2 m/sec from south.

25 bias

-- JohnHill - 12 May 2007
Topic revision: r5 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback